OpenStudy (anonymous):

At the candy store, a chocolate bar costs $c$ dollars and a vanilla bar costs $2$ dollars more than a chocolate bar. Jamie buys a chocolate bar and three vanilla bars, and Kevin buys five chocolate bars. How much money, in total, do Jamie and Kevin spend at the candy store in terms of $c$?

OpenStudy (anonymous):

Given, the CP of chocolate bar is$x Given, the CP of vanila bar is $x + $2(or simplly x + 2) Np. of chocolate Jamie brought = 1 Np. of chocolate Jamie brought = 3 Therefore,CP of Total no. of ice cream jamie brought(I1) = x + 3(x + 2) = 4x + 6 No. of chocolates Kevin brought = 5 therefore , CP of Total no. of ice cream kevin brought (II)= 5x SO total monney spent = (I) + (II) = 4x + 6 + 5x = 9x +6 @jenise